# FreeTacMan: Robot-free Visuo-Tactile Data Collection System for Contact-rich Manipulation

### v1

@@ 3.3 Tactile Pretraining and Policy Learning
Tactile Representation Learning. Our wearable system yields large-scale, contact-rich visuo-tactile trajectories, enabling a high-fidelity dataset for representation learning. Although tactile outputs resemble 2D images, applying a vision encoder pretrained on RGB data often produces suboptimal features due to the domain gap in appearance and semantics [17, 58]. We adopt a CLIP-style [59] contrastive pretraining procedure to bridge the domain gap.
@@
Both the visual encoder $f_{v}$ and tactile encoder $f_{t}$ share a ResNet backbone initialized from the same checkpoint. $f_{v}$ remains frozen during pretraining, while $f_{t}$ is finetuned. Each encoder is followed by a projection head: $g_{v}$ for vision, and $g_{t}$ for tactile, where $g_{t}$ first concatenates the tactile features with the normalized 7-DOF joint position vector $\mathbf{q}_{i}$ to inject robot joint state as global context. At each timestep $i$, we compute the normalized embeddings $\mathbf{v}_{i}$ and $\mathbf{t}_{i}$. We designate $\mathbf{v}_{i}$ as the primary positive for $\mathbf{t}_{i}$ to align tactile features with the current visual features. However, relying solely on time-aligned contrastive loss neglects temporal dynamics, leading to embeddings that vary abruptly and fail to capture evolving contact patterns. To enforce temporal awareness, a secondary positive $\mathbf{v}_{i+1}$ drawn from the next timestep is introduced. All other entries from negatives in a fixed-size memory bank $\mathcal{{M}}$. We train $f_{t}$, $g_{v}$ and $g_{t}$ by minimizing the following contrastive loss:
$\mathcal{L}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp\bigl{(}\mathbf{v}_{i}^{\top}\mathbf{t}_{i}/\tau\bigr{)}+\exp\bigl{(}\mathbf{v}_{i+1}^{\top}\mathbf{t}_{i}/\tau\bigr{)}}{\exp\bigl{(}\mathbf{v}_{i}^{\top}\mathbf{t}_{i}/\tau\bigr{)}+\exp\bigl{(}\mathbf{v}_{i+1}^{\top}\mathbf{t}_{i}/\tau\bigr{)}+\sum_{j\in\mathcal{N}_{i}}\exp\bigl{(}\mathbf{v}_{j}^{\top}\mathbf{t}_{i}/\tau\bigr{)}},$ (1)
@@
where $B$ indicates batch size, $\tau$ is a learned temperature parameter and $\mathcal{N}_{i}$ indexes the negatives. Please refer to Supplementary Appendix C for pretraining details.
@@
Visuo-Tactile Action Chunking Transformer. We employ the pretrained tactile encoder to extract tactile representations. The vision and tactile embeddings are then concatenated and input into the action chunking transformer (ACT) [11], which is trained to predict absolute joint positions.
4 Experiments
We design experiments to answer three key questions: Q1. Can demonstrations be collected efficiently and accurately using FreeTacMan compared to previous setups? Q2. Is tactile feedback improving the policy in contact-rich tasks, and how does the policy dynamically integrate visual and tactile information during the execution? Q3. What is the benefit of tactile encoder pretraining? How does the temporal-aware CLIP-style contrastive pretraining improve the policy learning?
4.1 Experimental Setup
@@
Figure 3: Human demonstrations and policy rollouts on our tasks. The top row shows the goal trajectory for each task, while the middle row demonstrates successful rollouts with tactile feedback. The bottom row showcases the typical failure mode of each task when tactile input is absent.
@@
To answer the questions above, we evaluate the effectiveness of FreeTacMan system and the quality of the collected visuo-tactile demonstration through a diverse set of contact-rich manipulation tasks: 1) Fragile Cup Manipulation. The robot grasps a small plastic cup and places it on a tray without causing damage. 2) USB Plugging. The robot needs to securely insert a pre-grasped USB plug into a socket. 3) Texture Classification. The robot should grasp and identify one of two cylindrical objects with distinct textures and sort it into the correct bin. 4) Stamp Pressing. The robot presses a stamp onto paper to produce a clear imprint. 5) Calligraphy Writing. The robot traces the digit "5" with a calligraphy brush. 6) Toothpaste Extrusion. The robot squeezes a tube to deposit a precise amount of toothpaste. 7) Tissue Grasping. The robot lifts a thin tissue from a bag. 8) Potato Chip Grasping. The robot picks up a fragile potato chip and places it onto a plate. These tasks cover a wide range of manipulation challenges such as force control, slip detection, and in-hand object pose detection. Five representative tasks are depicted in Fig. 3.
4.2 User Study on Data Collection System
@@
Figure 4: User Study on data collection. (a-c) FreeTacMan outperforms ALOHA [27] and UMI [30] in terms of completion rate, collection efficiency, and the CPUT score per task. (d) Average performance of all tasks across three evaluation criteria. (e) FreeTacMan excels at user experience evaluation as well, such as control accuracy, ease of collection procedure, and stability.
@@
Procedure. To answer (Q1), we evaluate the usability of FreeTacMan through a user study including 12 human volunteers, with varying levels of experience in data collection. They collect demonstrations for all 8 tasks. In addition to FreeTacMan, users collect demonstrations using two typical data collection methods: primary–replica-based teleoperation (i.e., ALOHA [27]) and handheld devices (i.e., UMI [30]). To minimize potential biases, no device-specific instructions are provided and each participant conducts three trials with each device for each task. There is no time limit for data collection, but they are instructed to solve tasks as best they can while avoiding collisions and damage. If a task fails, they will continue from where it is interrupted, and the failure will be recorded.
@@
Demonstrations could be collected efficiently and accurately using FreeTacMan compared to previous setups (Q1). The results of user study are shown in Fig. 4, where FreeTacMan consistently yields the top completion rate and efficiency. CPUT score in Fig. 4(c) highlights the overall advantage of FreeTacMan, gaining 6.58$\times$ higher performance than teleoperation and 7.68$\times$ higher than UMI.
@@
For simple tasks which do not require accurate force sensing or control, such as texture classification and tissue grasping, our system and UMI achieve comparable completion rates, while ALOHA performs slightly inferior. In terms of completion time, we have at least a 2.64×2.64$\times$2.64 × advantage over ALOHA and a 1.35×1.35$\times$1.35 × advantage over UMI. In more complex tasks, such as fragile cup manipulation and potato chip grasping, where force feedback is necessary to ensure successful grasping and prevent excessive force, primary-replica teleoperation leads to damage-related failures. This suggests that the handheld method provides better force feedback for grasping delicate objects.
@@
For tasks requiring dynamic force control, such as USB plugging, stamp pressing, and toothpaste extrusion, ALOHA relies on brute force, barely completing them at the risk of causing damage, while UMI often fails due to the lack of slippage detection. In contrast, FreeTacMan combines both precise force perception and control, resulting in superior accuracy and safety. For high-precision trajectory control (e.g., calligraphy writing), ALOHA struggles significantly—users often need to manually assist the teaching arm with their other hand, resulting in the longest completion time. While UMI can complete the task, the trajectory smoothness remains inferior to that of FreeTacMan.
@@
Fig. 4(e) summarizes the user experience evaluation. Stability, comfort, ease of collection and entry are assessed via questionnaires and normalized. Stability and damage represent failures of object drops and damage. The results show that FreeTacMan is the most user-friendly and reliable data collection system among the three approaches.
@@
Table 3: Policy success rates (%) across contact-rich tasks. The visuo-tactile information, together with the pretraining strategy, greatly helps imitation learning for the contact-rich tasks.
Method Fragile Cup USB Plug Texture Cls. Stamp Press Calligraphy Avg.
ACT [11] (Vision-only) 35 0 20 20 30 21
Ours (+ Tactile w/o Pretraining) 75 10 70 55 65 55
Ours (+ Pretraining) 80 20 90 85 80 71
4.3 Validation on Imitation Learning
We collect data and train policies with tasks in Fig. 3 and evaluate each task over 20 trials.
@@ 4.3 Validation on Imitation Learning
  • ACT [11] (Vision-only): The baseline ACT uses RGB images from the wrist camera as input only.
@@
  • Ours (+ tactile w/o pretraining): An extended ACT model taking both visual and tactile observations, which are separately encoded by identical backbones without pretraining.
@@
  • Ours (+ pretraining): Our full model, where the tactile encoder is pretrained with a multi-positive contrastive objective incorporating both primary and secondary positives.
@@
Integration of tactile feedback results in significant improvement on policy performance, confirming its dynamic value in contact-rich tasks (Q2). The vision-only baseline ACT achieves low performance across all tasks, with an average success rate of 21%. Without tactile feedback, the robot struggles in precise manipulation tasks such as USB plugging and force-sensitive tasks like fragile cup. When tactile feedback is incorporated naively, i.e., without pre-training, performance improves significantly, with the average success rate increasing to 55%. Tasks such as fragile cup handling (75%) and texture classification (70%) gain a remarkable boost over the vision-only baseline. This demonstrates the utility of tactile sensing in tasks that involve contact-rich interactions and visual cues alone are insufficient to distinguish subtle features. However, the performance in tasks like USB plug (10%) and stamp press (55%) still lags behind, indicating room for further refinement.
@@
Temporal-aware pretraining improves the policy performance by aligning visual and tactile embeddings and incorporating temporal dynamics (Q3). Incorporating time-aware visual-tactile pairs in pretraining leads to a notable performance boost, with the average success rate increasing to 71%. As shown in the last line in Table 3, the improvement comes from tasks that demand fine-grained control and the ability to track contact dynamics, such as stamp pressing (85%) and calligraphy (80%). Meaningful gain is witnessed in USB plug (20%). These results show that our approach aligns tactile and visual embeddings while incorporating spatial and temporal context.
5 Conclusion
@@
We present FreeTacMan, a human-centric and robot-free data collection system with in-situ visuo-tactile feedback and recording. A high-quality visuo-tactile sensor is developed, and an in-situ and modular gripper is designed for rapid adaptation. Experimental results demonstrate that the proposed system outperforms existing methods in multiple aspects, including data collection efficiency, control accuracy, and human-machine interaction experience. Through policy validation, the effectiveness of the collected data and the importance of visuo-tactile pretraining are further confirmed.
Limitations
While FreeTacMan has demonstrated efficacy across a range of challenging tasks, a few limitations remain. First, our system currently depends on an external base station for localization. We plan to eliminate this dependency in the future by introducing high-precision visual algorithms, with a continuous focus on reducing hardware reliance and further enhancing system performance. Besides, we have not released a large-scale dataset in this work. However, our data collection system has paved the way to rapidly gather large-scale multi-modal data thanks to its accuracy and efficiency, and we would prioritize this goal in our future work. In addition, we employ a basic policy model, ACT, for downstream validation, but the integration of the visuo-tactile information has not yet been fully investigated. It is worth exploring advanced visuo-tactile sensor fusion algorithms.
Acknowledgments

### v2

@@ III-C Tactile Pretraining and Policy Learning
Figure 3: Tactile pretraining and policy learning pipeline. (a) A tactile encoder is pretrained using the self-collected dataset. (b) The pretrained tactile encoder is integrated into an ACT-based policy for downstream tasks such as USB insertion.
@@
Tactile representation learning. Our wearable system yields high-precision, contact-rich visuo-tactile trajectories, enabling a high-fidelity dataset for representation learning. Although tactile outputs resemble 2D images, applying a vision encoder pretrained on RGB data often produces suboptimal features due to the domain gap in appearance and semantics [9, 42]. We adopt a CLIP-style [43] contrastive pretraining procedure to bridge the domain gap.
@@
As illustrated in Fig.˜3(a), both the visual encoder $f_{v}$ and tactile encoder $f_{t}$ share a ResNet backbone initialized from the same checkpoint. $f_{v}$ remains frozen during pretraining, while $f_{t}$ is finetuned. Each encoder is followed by a projection head: $g_{v}$ for vision, and $g_{t}$ for tactile, where $g_{t}$ first concatenates the tactile features with the normalized 7-DOF joint position vector $\mathbf{q}_{i}$ to inject robot joint state as global context. At each timestep $i$, we compute the normalized embeddings $\mathbf{v}_{i}$ and $\mathbf{t}_{i}$. We designate $\mathbf{v}_{i}$ as the primary positive for $\mathbf{t}_{i}$ to align tactile features with the current visual features. However, relying solely on time-aligned contrastive loss neglects temporal dynamics, leading to embeddings that vary abruptly and fail to capture evolving contact patterns. To enforce temporal awareness, a secondary positive $\mathbf{v}_{i+1}$ drawn from the next timestep is introduced. All other entries from negatives in a fixed-size memory bank $\mathcal{{M}}$. We train $f_{t}$, $g_{v}$ and $g_{t}$ by minimizing the following contrastive loss:
$L=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{e^{\mathbf{v}_{i}^{\top}\mathbf{t}_{i}/\tau}+e^{\mathbf{v}_{i+1}^{\top}\mathbf{t}_{i}/\tau}}{e^{\mathbf{v}_{i}^{\top}\mathbf{t}_{i}/\tau}+e^{\mathbf{v}_{i+1}^{\top}\mathbf{t}_{i}/\tau}+\sum_{j\in\mathcal{N}_{i}}e^{\mathbf{v}_{j}^{\top}\mathbf{t}_{i}/\tau}},$ (1)
@@
where $B$ indicates batch size, $\tau$ is a learned temperature parameter and $\mathcal{N}_{i}$ indexes the negatives.
@@
Visuo-tactile action chunking transformer. We employ the pretrained tactile encoder to extract tactile representations. As shown in Fig.˜3(b), vision and tactile embeddings are then concatenated and input into the action chunking transformer (ACT) [26], which is trained to predict joint positions.
IV Experiments
We design experiments to answer three key questions:
Q1. Can demonstrations be collected efficiently and accurately using FreeTacMan compared to previous setups?
Q2. Is the in-situ tactile information in FreeTacMan dataset effective for contact-rich tasks policy learning?
Q3. How dose the tactile encoder pretrained on self-collected visuo-tactile data improve policy learning?
IV-A Dataset
Figure 4: The FreeTacMan dataset. (a) Representative examples illustrating the diversity in both task complexity and tactile context. (b) The dataset covers 50 tasks and features a large-scale collection of data, including more than 10k trajectories which contains over 3000k visuo-tactile pairs. (c) The dataset enables diverse fundamental tactile capabilities.
Enabled by the efficient, precise, and fidelity tactile data collection system, we curate a diverse dataset of manipulation tasks spanning vision, touch, and proprioception modalities, as illustrated in Fig.˜4. The dataset spans 50 tasks, comprising more than 10k manipulation trajectories which contains over 3000k visuo-tactile image pairs. Collected with high-consistent and high-resolution visuo-tactile sensors, it enables large-scale tactile pretraining and multimodal policy learning. Furthermore, the dataset covers diverse fundamental tactile capabilities, as shown in Figure˜4(c). To ensure dataset quality, every trajectory is replayed in our validation process, providing a robust filter that guarantees the reliability of the data for downstream robot learning.
IV-B Experimental Setup
@@
Figure 5: Human demonstrations and policy rollouts. The top row shows goal trajectories, the middle row demonstrates successful rollouts with tactile feedback, and the bottom row showcases typical failure modes when without tactile input.
@@
To answer the questions above, we evaluate the effectiveness of FreeTacMan system and the quality of the collected dataset through a diverse set of contact-rich manipulation tasks, shown in Figure˜5. 1) Fragile cup. The robot grasps a fragile cup without causing damage and places it stably on a tray. 2) Calligraphy. The robot traces the digit "5" with a calligraphy brush. 3) Texture classification. The robot grasps and identifies one of two cylindrical objects with distinct textures and sorts it into the correct bin. 4) Stamp press. The robot presses a stamp onto paper to produce a clear imprint. 5) USB plug. The robot needs to securely insert a pre-grasped USB plug into a socket (error <1mm<1mm). These tasks collectively represent the four core tactile capabilities of our dataset: force control (fragile cup), hybrid force-position control (USB plug, stamp press), high-acuity tactile perception (texture classification) and dynamic response (calligraphy).
IV-C User Study on Data Collection System
@@
Figure 6: User Study on data collection. (a-c) FreeTacMan outperforms ALOHA [44] and UMI [15] in terms of completion rate, collection efficiency, and the CPUT score per task. (d) FreeTacMan excels at user experience evaluation as well.
Figure 7: Trajectory visualization. We test FreeTacMan on a variety of contact-rich tasks. Videos are available on our website.
@@
Procedure. We evaluate the usability of FreeTacMan through a user study including 12 volunteers, with varying experience in data collection. Besides FreeTacMan, users collect demonstrations using two typical methods: primary–replica-based teleoperation (i.e., ALOHA [44]) and handheld devices (i.e., UMI [15]). To minimize biases, no device-specific instructions are provided and each participant conducts three trials with each device for each task. The participants are instructed to solve tasks as best they can while avoiding collisions and damage. If a task fails, they will continue from where it is interrupted, and the failure will be recorded.
Metrics. We record the task success/failure mode, completion time, and any instances of slippage or damage. Three metrics are adopted to quantify the capability of a particular data collection approach, namely completion_rate (fully completed tasks as a percentage of those initiated), collection efficiency (the inverse of data collection time), and an overall score, Completion per Unit Time (CPUT), defined as $\texttt{completion\_rate}\times\texttt{efficiency}$.
@@
For simple tasks like texture classification that don’t require accurate force sensing or control, our system and UMI achieve comparable completion rates, while ALOHA performs slightly inferior. In completion time, we have a $2.65\times$ advantage over ALOHA and a $1.22\times$ advantage over UMI. In more complex tasks, such as fragile cup manipulation, where force feedback is necessary to ensure successful grasping and prevent excessive force, primary-replica teleoperation leads to damage-related failures. This suggests that the handheld method provides better force feedback for delicate objects.
@@
For tasks requiring dynamic force control, such as USB plug and stamp press, ALOHA relies on brute force, barely completing them at the risk of causing damage, while UMI often fails due to the lack of slippage detection. In contrast, FreeTacMan combines both precise force perception and control, resulting in superior accuracy and safety. For high-precision trajectory control (e.g., calligraphy), ALOHA struggles significantly—users often need to manually assist the teaching arm with their other hand, resulting in the longest completion time. While UMI can complete the task, the trajectory smoothness remains inferior to that of FreeTacMan.
@@
Fig.˜6(d) summarizes the user experience evaluation. Stability, comfort, ease of collection and entry are assessed via questionnaires and normalized. Stability and damage represent failures of object drops and damage. The results show that FreeTacMan is the most user-friendly and reliable data collection system among the three approaches.
@@
TABLE II: Policy success rates (%) across tasks. Tactile input and pretraining significantly boost imitation learning.
Method
Fragile
Cup
Calligraphy
Texture
Cls.
Stamp
Press
USB
Plug
Avg.
ACT [26]
(Vision-only)
35 20 20 30 0 21
Ours (+ Tactile
w/o Pretrain)
75 70 55 65 10 55
Ours
(+ Pretrain)
80 90 85 80 20 71
IV-D Validation on Imitation Learning
Each task in Fig.˜7 is trained and evaluated over 20 trials.
@@ IV-D Validation on Imitation Learning
  • ACT [26] (Vision-only): The original ACT uses RGB images from the wrist camera as input only.
@@
  • Ours (+ Tactile w/o Pretrain): An extended ACT model taking both visual and tactile observations, which are separately encoded by identical backbones without pretraining.
@@
  • Ours (+ Pretrain): Our full model, where the tactile encoder is pretrained with a multi-positive contrastive objective incorporating both primary and secondary positives.
@@
Integration of tactile feedback results in significant improvement on policy performance (Q2). As illustrated in Table˜II, the vision-only baseline achieves low performance across all tasks, with an average success rate of 21%. Without tactile feedback, the robot struggles in tasks requiring fine tactile perception, such as USB plug, calligraphy, and texture classification. When tactile feedback is incorporated naively, i.e., without pretraining, performance improves substantially, with the success rate increasing to 55%. Significant gains are observed in tasks such as fragile cup manipulation (75%) and calligraphy (70%). This highlights the utility of tactile sensing in contact-rich tasks where visual cues alone are insufficient to distinguish subtle features. However, the performance in tasks like USB plug (10%) and texture classification (55%) still lags behind, indicating room for further refinement.
Tactile inputs enables excellent performance even on unseen objects (Q2). In the texture classification task, we select a texture object with an out-of-distribution color (red) for testing. As shown in Table˜III, tactile input achieves 55% accuracy on unseen objects, matching performance on training objects and significantly exceeding the 15% vision-only baseline. Pretraining further increases accuracy to 70%, demonstrating robust generalization.
TABLE III: Comparison in unseen object. Tactile input enables generalization to unseen objects.
Vision -only Ours (+ Tactile w/o Pretrain) Ours (+ Pretrain)
training object 20% 55% 85%
unseen object 15% 55% 70%
@@
Temporal-aware pretraining improves the performance by aligning visual and tactile embeddings (Q3). Incorporating both time-aligned and time-adjacent visual-tactile pairs in CLIP pretraining leads to a notable performance boost, with the average success rate increasing to 71%. As shown in the last line in Table˜II, the improvement comes from tasks that demand fine-grained control and the ability to track contact dynamics, such as calligraphy (90%) and stamp (80%). Meaningful gain is witnessed in USB plug (20%). These results show that our approach aligns tactile and visual embeddings while incorporating spatial and temporal context.
Pretraining demonstrates substantial benefits under data-scarce conditions (Q3). As shown in Fig.˜8, ACT policy with pretrained encoder achieves superior performance with only 50 task-specific demonstrations—55% for cup manipulation, 50% for calligraphy and 60% for texture classification—compared to the visual-only baseline (35%, 20%, and 20%, respectively). With 100 episodes, its performance in cup and stamp tasks (60% each) nearly matches the non-pretrained tactile model (70% and 65%). We further analyze performance discrepancies across task types. Perception-centric tasks like texture classification benefit from the generalized features of the pre-training, while action-oriented tasks such as cup manipulation and calligraphy require not only robust features but also sufficient task-specific demonstrations to learn effective action strategies.
Figure 8: Ablation across training episodes Our pretrained policy, even with small data, outperforms vision-only policy trained with large amounts of data. Moreover, it approaches the performance of a large-data tactile-only policy.
Pretraining on the FreeTacMan dataset provides robustness against variations among visuo-tactile sensors (Q3). While our main experiments are conducted in an in-domain setting, where training and evaluation use different sensor batches with high consistency, we also evaluate a more challenging out-of-domain scenario by introducing substantial variations across visuo-tactile sensors. As shown in Table˜IV, pretrained model generalizes well across two distinct sensors, which differ in marker angle (0° vs. 45°) and RGB lighting environment (side vs. bottom). The in-domain and out-of-domain performance remain close on the texture classification task (0.85 vs. 0.75) and identical on the fragile cup task (both 0.8). In contrast, the non-pretrained counterpart performs poorly, particularly in phases requiring fine-grained tactile perception: a drop from 80% to 40% in the tactile-based classification phase, and a catastrophic drop from 70% to 10% in the landing phase (which relies on detecting contact with table before releasing). These results validate the effectiveness of tactile pretraining for cross-sensor generalization.
TABLE IV: Cross-sensor generalization results. Visuo-tactile pretraining helps generalization to other visuo-tactile sensor setups. ID: in-domain. OOD: out-of-domain.
Task Phase Ours (+ Tactile w/o Pretrain) Ours (+ Pretrain)
ID Sensor OOD Sensor ID Sensor OOD Sensor
Texture Cls. Grasp 75% 80% 85% 85%
Classify 55% 40% 85% 75%
Whole Task 55% 40% 85% 75%
Fragile Cup Pick 80% 70% 80% 80%
Land 75% 10% 80% 80%
Whole Task 75% 10% 80% 80%
V Conclusion
@@
We present FreeTacMan, a human-centric and robot-free data collection system with in-situ visuo-tactile feedback and recording. An in-situ and modular gripper with visuo-tactile sensors enables rapid adaptation, and a large-scale high-precision dataset is collected to support contact-rich manipulation policy learning. Experimental results demonstrate that the proposed system outperforms existing methods in multiple aspects, including data collection efficiency, control accuracy, and human-machine interaction experience. Through policy validation, the effectiveness of the collected data and the importance of visuo-tactile pretraining are further confirmed.
Limitation and future work. While FreeTacMan has demonstrated efficacy across a range of challenging tasks, a few limitations remain. Our system currently depends on an external base station for submillimeter-level localization. To eliminate this dependency, we will develop high-precision visual algorithms for collecting precise visuo-tactile data in the wild. We also plan to extend our data collection system and dataset to multi-fingered dexterous hands, enabling the study of more complex and fine-grained manipulation scenarios.
VI Acknowledgments